A hydraulic press has a small piston of diameter 2 cm. If a force of 60 N is applied to the smaller

piston, a force of 3000 N is exerted on the larger piston. What is the diameter of the larger piston?


1
Expert's answer
2021-05-09T15:13:20-0400

By pascal's law we can say that pressure inside , the external static pressure applied on a confined liquid is distributed or transmitted evenly through out the liquid in all direction.


P1 = P2

=

"\dfrac{F_1}{A_1}" "= \dfrac{F_2}{A_2}"


=

"\dfrac{60}{{\pi}r_1^{2}}=" "\dfrac{3000}{{\pi}r_2^{2}}" .....1)


where r1 , r2 are respectively radius of smaller and larger piston .


r1 = 1 cm

Putting the values in equation 1)


"r_2^{2}" ="\dfrac{3000}{60}=50"


"r_2=7.07\ cm."